OpenStudy (anonymous):

Find f^-1(x). f(x) = 3 square root of x - 4

OpenStudy (anonymous):

\[f(x) = \sqrt[3]{x-4}\]

OpenStudy (anonymous):

y = cube root of x -4 y^3 = x - 4 x = y^3 + 4 f^-1 (x) = x^3 +4